What is the maturity value of a $100,000, 5%, six-month note? O A. $97,500 OB. $100,000 OC. $102,500 OD. $105,000
Answer #1
FV= PV*(1+r)^n
Where,
FV= Future Value
PV = Present Value
r = Interest rate i.e. 5*6/12 =2.5%
n= Number of years
= $100000*( 1+0.025)^1
=100000*1.025
= $102500
Correct Option: C